# Double strangeness pentaquarks and other exotic hadrons in the $\Xi_{b}\to\Xi\ J/\Psi\ \phi$ decay

### v1

@@ IV.1 Dalitz Plot
As it was mentioned earlier in Ref. Oset:2024fbk the authors suggest the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decays in order to observe the $P_{css}$ with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$ generated via pseudoscalar-baryon interaction. In this study the authors used a model from ref. Roca:2024nsi , i.e. with narrow $P_{css}$ pentaquarks.
In this chapter we will study the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay, as a possible reaction to look for a double strangeness pentaquark via vector-baryon channel. And first of all, we will look at Dalitz plots, i.e. at the two-body invariant mass distributions of the three possible pairs of particles $(J/\psi\phi,~{}J/\psi\Xi,~{}\phi\Xi)$, what will make clear that there exists a good possibility to detect exotic hadrons and not only pentaquarks $P_{css}$.
From the relativistic kinematics of a three body decay the Dalitz plot can be generated, which represents the available invariant mass for the different pairs of particles. So if we explore the Dalitz plots for the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ reaction (Fig. 15) we can see whether the different exotic resonances lie inside the kinematically allowed region.
@@
On the first two Dalitz plots (Fig. 15 (a) and (b)) the dashed vertical line corresponds to the mass predicted for the $S=-2$ pentaquark and the shadowed zone shows the variations of its mass of ±$\simeq\pm 50~{}{\rm MeV}$ due to theoretical uncertainties. We note that, if the pentaquark with double strangeness, discussed later in this work, appears inside the allowed kinematical region, then it may contribute with an important peak in the corresponding invariant mass distribution.
Other examples of exotic states are the $X$ meson resonances, which can be observed in the observed in ${J/\psi\phi}$ invariant mass. Dalitz plots (Fig. 15 (b) and (c)) show that five $X$ resonances lie inside the permitted kinematical region, namely the $X(4140)$, $X(4160)$, $X(4274)$, $X(4550)$ and the $X(4700)$. In spite of the fact that all five may be detectable, we will not consider the last tree states since there is at the moment no theoretical model that describes them.
@@
There is still a discussion of the nature and properties of the $X(4140)$ and $X(4160)$ states. In 2008, the Belle collaboration reported the existence of the $X(4160)$ Belle:2007woe in the $e^{-}e^{+}\to J/\psi D^{*}\bar{D}^{*}$ reaction and, during the following years, some groups reported the existence of a narrow $X(4140)$ with a width around $19~{}{\rm MeV}$ CDF:2009jgo ; CDF:2011pep ; LHCb:2012wyi ; CMS:2013jru ; D0:2013jvp ; D0:2015nxw . Despite of these results, a more recent measurement of the $B^{+}\to J/\psi\phi K^{+}$ decay from the LHCb collaboration LHCb:2016axx obtained a width of $83~{}{\rm MeV}$ for the $X(4140)$, which is pretty large compared with the previously studies. In that work other states that couple to $J/\psi\phi$ were also reported such as the $X(4274)$, $X(4550)$ and the $X(4700)$, but the $X(4160)$, which is the state that was previously seen in 2008, was not observed. It might be that the reason of this large width for $X(4140)$ is related to the fact that the two neighbouring states (the narrow $X(4140)$ state plus a wider $X(4160)$ resonance) were fitted together. In this work we will study the interplay between the $X(4140)$ and the $X(4160)$ resonances and will try to find an observable that permits us to discriminate whether the truly nature of this state is only one wide $X(4140)$ resonance or it is the combination of a narrow $X(4140)$ state plus a $X(4160)$ one.
@@
Among the theoretical works studying these $X$ states, there were some groups trying to identify the $X(4140)$ as a molecular state of a $D_{s}^{*}\bar{D}_{s}^{*}$ pair with the quantum numbers $0^{++}$ and $2^{++}$ Liu:2009ei ; Branz:2009yt ; Chen:2015fdn , but these studies did not take into account the light meson states, producing a small width for the state, hence it was associated to the $X(4140)$. Later, in Ref. Molina:2009ct , the contribution of the light meson channels were included, generating a $2^{++}$ $X$ state at $4169~{}{\rm MeV}$ with a width of $139~{}{\rm MeV}$ that strongly couples to $D_{s}^{*}\bar{D}_{s}^{*}$ and it was associated to the $X(4160)$. In the end the quantum numbers of $X(4140)$ were measured, obtaining that these were $1^{++}$ PDG , therefore the models that predict this state as a $D_{s}^{*}\bar{D}_{s}^{*}$ molecule are no longer supported.
@@ IV.2 Primary decay
Studying the $\Xi_{b}\to\Xi X\to\Xi~{}J/\psi~{}\phi$ decay we, apart from the search for the double strangeness pentaquark, have also a chance to study the interplay between the $X(4140)$ and $X(4160)$ resonances. Similar to what was done in Refs. Magas:2020zuo ; Wang:2017mrt , we use two independent primary decay mechanisms, one for $X(4160)$ and another for $X(4140)$. Since for the $X(4140)$ resonance we do not have a physical model to generate it, we introduce it as a Breit-Wigner whose parameters are fitted to experimental data. In contrast, for the $X(4160)$ we do have a molecular model which predicts rather well its mass and width (Ref. Molina:2009ct ), and also predicts its strong coupling to $D_{s}^{*}\bar{D}_{s}^{*}$ channel.
@@
The dominant process for the $\Xi_{b}$ decay at the quark level proceeds via the external emission mechanism depicted in Fig. 16. Since the final state involves $J/\psi\,\phi$ pairs, the $D_{s}^{*}\,\bar{D}_{s}^{*}$ lines have to be connected to form a loop, as shown in Fig. 17, allowing multiple coupled channel interactions and generating dynamically the $X(4160)$, which then decays into $J/\psi$ and $\phi$. According to ref. Molina:2009ct the $X(4160)$ couples with maximum strength to the $D_{s}^{*}\,\bar{D}_{s}^{*}$ channel, but its coupling to the $J/\psi\,\phi$ channel is also significant, what also increase the dominance of the process depicted in Fig. 17. On the other hand, if we draw a diagram which would be responsible even for a direct $J/\psi\,\phi\Xi$ decay, at the microscopic quark level this reaction proceeds via the internal conversion process, shown in Fig. 18, which is strongly penalized by color factors color_fac and therefore it will be neglected in this work, similar to what was done in Ref. Magas:2020zuo ; Wang:2017mrt .
@@
The $A$ on the right side of the equation represents the strength of the $\Xi_{b}\to\Xi D_{s}^{*}~{}\bar{D}_{s}^{*}$ weak decay, shown in Fig. 16, which we are not going to study in details. This value can be taken as a constant due to the limited range of energies involved in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay as it is argued in Ref. Feijoo:2015cca . The $(\vec{\epsilon}_{J/\psi}\times\vec{\epsilon}_{\phi})\cdot\vec{P}_{\Xi}$ factor denotes the P-wave operator which is the minimum partial wave we need in the weak vertex to conserve the angular momentum, since the spins of $\Xi_{b}$ and $\Xi$ are $J=1/2$ while for the $X(4160)$ is $J=2$. Here $\vec{\epsilon}_{i}$ represents the polarization of the vector mesons in the $J/\psi~{}\phi$ rest frame and $\vec{P}_{\Xi}$ the tree-momentum of the $\Xi$ in the same frame. The factor $G_{D_{s}^{*}\bar{D}_{s}^{*}}$ indicates the contribution of the $D_{s}^{*}\bar{D}_{s}^{*}$ loop shown in Fig. 17 and the $T_{D_{s}^{*}\bar{D}_{s}^{*},J/\psi\phi}$ term is the coupled-channel unitarized amplitude for the $D_{s}^{*}\bar{D}_{s}^{*}\to J/\psi~{}\phi$ process. Note that we divide it by the corresponding couplings constants. This is done for a proper comparison with the $X(4140)$ term.
Now we will discuss the alternative way of our reaction via production of $X(4140)$, as it is shown in Fig. 19. The associated amplitude is
@@
where the $X(4140)$ is parameterised with a Breit-Wigner, where $M_{J/\psi\phi}$ is the invariant mass of the $J/\psi\phi$ system and $M_{X(4140)}$ and $\Gamma_{X(4140)}$ are the mass and width of $X(4140)$ resonance, respectively. A new constant, connected to the strength of the $\Xi_{b}\to\Xi~{}X(4140)$ weak decay, is $B$ and, since we want it to have the same units as $A$, we introduce a $2M_{X(4140)}$ factor in the denominator. In this process we do not need to have a P-vertex since the quantum number of the $X(4140)$ is $J=1$. Therefore the minimum partial wave needed to conserve the angular momentum is $L=0$.
@@
We will not go into more detailed calculations, since the main aim of this work is to point out the possibility to use the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay to study already discovered and potentially new exotic states and to encourage to the experimental collaborations to explore this reaction.
@@ IV.3 Final state interaction
Once we discussed the generation of the $\Xi~{}J/\psi~{}\phi$ final state in $\Xi_{b}$ decay we want to focus on studying the final state interaction in the $\Xi~{}J/\psi$ and $\Xi~{}\phi$ channels. For the case of the final state interaction of $\Xi~{}J/\psi$ pair the $\Xi$ and $J/\psi$ legs from Fig. 17 and 19 have to be connected, forming a loop, which can dynamically generate the $S=-2$ pentaquark. The $\Xi~{}\phi$ final state interaction can produce excited $\Xi$ states and has to be studied in a similar way.
@@
The amplitude associated to the process with the final state interaction of the $J/\psi\Xi$ pair driven by the $\Xi_{b}$ decay via $X(4160)$ resonance is:
@@
which is associated to the diagram shown in Fig. 20(a), where the superscript denotes the final state interaction channel. This expression is derived in Appendix A, where the expression for the integral $I^{J/\psi\Xi}_{X_{4160}}$ is also given. We would like to remind the reader that for the reaction via $X(4160)$ the weak decay vertex has to be a P-wave one. Therefore the operator assigned to the the $\Xi_{b}\to D^{*}_{s}~{}\bar{D}^{*}_{s}~{}\Xi$ is proportional to the momentum of $\Xi$ and the loop integral (Fig. 20(a)) becomes a tree-vector. We evaluate this loop integral in the $J/\psi~{}\Xi$ rest frame, where the loop integral becomes proportional to $\frac{(\vec{P}_{\Xi}-\vec{P}_{\phi})}{2}$, with the scalar coefficient $I^{J/\psi\Xi}_{X_{4160}}$. The term $T_{J/\psi\Xi,J/\psi\Xi}$ captures the final state interaction contribution from $J/\psi~{}\Xi$, which in this work we will model using a Breit-Wigner form:
@@
were $g_{J/\psi~{}\Xi}$ is the coupling of the $S=-2$ pentaquark to the $J/\psi~{}\Xi$ channel. $M_{J/\psi~{}\Xi}$ denotes the invariant mass of the $J/\psi~{}\Xi$ system, and $M$ and $\Gamma$ denote the mass and the width of the $S=-2$ pentaquark. The values of $g_{J/\psi~{}\Xi}$, $M$ and $\Gamma$ will be obtained from the molecular model of a VB interacting discussed in section II.
@@
where the $T_{\phi\Xi,\phi\Xi}$ amplitude is also taken as a Breit-Wigner form, as in Eq. (26), but in this case $M$ and $\Gamma$ are the experimental mass and width of the $\Xi(2500)$. Since we do not have any model which can give us the value of $g_{\phi\Xi}$, we will invert the problem and will study for which values of $g_{\phi\Xi}$ it is feasible to see this state using the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ reaction.
Now we go to the final state interaction in the presence of the $X(4140)$ resonance. The corresponding diagrams are given in Figs. 21(a) and 21(b) and the associated analytical expressions are
@@
These expressions are explained in more detail in Appendix A. The terms $I^{J/\psi\Xi}_{X_{4140}}$ and $I^{\phi\Xi}_{X_{4140}}$ are scalar loop integral analogous to those in Eqs. (25) and (27).
@@ IV.4 Full amplitude and decay rate
Now we proceed to combine the various terms to form the full invariant amplitude. Firstly we put together the $X(4140)$ terms into a ${\cal M}_{X_{4140}}$ amplitude, and similarly for $X(4160)$ terms,
@@
where we wrote explicitly the primary decay strengths $A$ and $B$. And secondly, we mix these terms obtaining the full amplitude denoted as ${\cal M}$,
@@
where the bar represents the sum over polarizations and $\beta=|B|^{2}/|A|^{2}$. Note that, since the weak decay goes in P-wave for the $X(4160)$ contribution and in S-wave for the $X(4140)$ contribution, the cross term in $\overline{{|\cal M}_{4160}+{\cal M}_{4140}|^{2}}$ cancels as these two partial waves are orthogonal and do not interfere. It is also important to comment that, although the the overall factor $|A|^{2}$ is not known, it is not relevant for the shape of the obtained distributions. The form of the distributions, and not its absolute value, will be our main observable. We are exploring if the position of the peak and its width can be measured in several model situations. So our final results will be given in arbitrary units, therefore we will set $|A|^{2}=1$ from here to the end of this work:
@@
The $\beta$ acts like a relative weight between the $X(4140)$ and $X(4160)$ contributions. In this work we will not calculate its value and we do not have any experimental result to compare it, so we take the value of $\beta$ from Ref. Wang:2017mrt to solve this issue. In this reference the authors study the interplay between $X(4160)$ and $X(4140)$ in the $B^{+}\to J/\psi~{}\phi~{}K^{+}$ decay. The mechanism involving the $X(4160)$ in this reaction is rather similar at the quark level to our diagram shown in Fig. 16, but without the spectator $s$ quark in the initial and final state. Hence the topology of the diagrams is similar also in the case of $X(4140)$. Therefore the $\beta$ value in our reaction may be similar as that in Ref. Wang:2017mrt , but we do not expect it to be the exactly the same, since the partial waves involved in their weak decay vertex are P-wave and D-wave for the $X(4140)$ and $X(4160)$ resonances respectively.
Finally, we explain how we deal with the sum over the polarizations appearing in Eq. (33). For the ${\cal M}_{X(4140)}$ amplitude this calculation is trivial and it only produces a constant, which can be reabsorbed into $\beta$, but for the ${\cal M}_{X(4160)}$ amplitude the calculation is more complicated adding cross term contributions. The full derivation is presented in Appendix B.
First of all, we introduce the following definitions:
@@
Finally, the double differential cross-section for the $\Xi_{b}\to J/\Psi\phi\Xi$ decay process reads rocamai :
@@
where $\mathcal{M}$ is given by eq. (33). As an example, we wrote in eq. (IV.4) cross section in function of $M_{\phi\Xi}$ and $M_{J/\psi\Xi}$ invariant masses, but we can write similar expression for any other combination of the two particle invariant masses of the $J/\Psi\phi\Xi$ final state (since out of three possible combination only two are independent).
Then, for example, fixing the invariant mass $M_{J/\psi\Xi}$, one can integrate over $M_{\phi\Xi}$ in order to obtain $d\Gamma/dM_{J/\psi\Xi}$. In this case, the limits are given by:

### v2

@@ IV.1 Dalitz Plot
As mentioned earlier, the authors of Ref. [27] suggest the measurement of the $\Xi_{b}^{0}\to\eta\eta_{c}\Xi^{0}$ and $\Omega_{b}^{-}\to K^{-}\eta_{c}\Xi^{0}$ decays in order to observe the $P_{css}$ state with $I(J^{P})=\frac{1}{2}(\frac{1}{2}^{-})$, generated from a pseudoscalar-baryon interaction, in the invariant mass of the $\eta_{c}\Xi^{0}$ pairs.
In the present work we investigate the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay as a possible process to look for a double strangeness pentaquark, coupling strongly to a vector meson-baryon pair, in the $J/\psi\Xi$ invariant mass. We will first inspect the Dalitz plots representing the regions of two-body invariant mass distributions of the three possible pairs of particles $(J/\psi\phi,~{}J/\psi\Xi,~{}\phi\Xi)$ to show that there exists a good possibility to detect not only pentaquarks $P_{css}$, but other exotic hadrons as well.
@@
In the first two Dalitz plots represented in Fig. 14 (panels (a) and (b)) the dashed vertical line corresponds to the mass of the $S=-2$ pentaquark predicted by our $VB$ model ($4633$ MeV) and the shadowed zone shows the variations of its mass of ±$\simeq\pm 50~{}{\rm MeV}$ accounting for theoretical uncertainties. It is clear that the invariant mass range of the $J/\psi\Xi$ pairs covers amply the region where our double strangeness pentaquark is predicted and it should therefore leave a signal in their spectrum.
@@
Other examples of exotic states are the $X$ meson resonances, which can be observed in the ${J/\psi\phi}$ invariant mass spectrum. The Dalitz plots (b) and (c) of Fig. 14 show that five $X$ resonances, the $X(4140)$, $X(4160)$, $X(4274)$, $X(4550)$ and the $X(4700)$, lie inside the permitted kinematical region. In spite of the fact that all five states may be detectable, we will only focus on the first two states, as they have been the object of some controversy. Indeed, there is still a discussion of the nature and properties of the $X(4140)$ and $X(4160)$ states. In 2008, the Belle collaboration reported the existence of the $X(4160)$ [49] in the $e^{-}e^{+}\to J/\psi D^{*}\bar{D}^{*}$ reaction and, during the following years, some groups reported the existence of a narrow $X(4140)$ with a width around $19~{}{\rm MeV}$ [50, 51, 52, 53, 54, 55]. Despite of these results, a more recent measurement of the $B^{+}\to J/\psi\phi K^{+}$ decay from the LHCb collaboration [56] obtained a width of $83~{}{\rm MeV}$ for the $X(4140)$, which is pretty large compared with the previously studies. In that work other states that couple to $J/\psi\phi$ were also reported such as the $X(4274)$, $X(4550)$ and the $X(4700)$, but the $X(4160)$, which is the state that was previously seen in 2008, was not observed. It might be that the reason of this large width for $X(4140)$ is related to the fact that the two neighboring states (the narrow $X(4140)$ state plus a wider $X(4160)$ resonance) were fitted together. In this work we will study the interplay between the $X(4140)$ and the $X(4160)$ resonances and will try to find an observable that permits us to discriminate whether the truly nature of this state is only one wide $X(4140)$ resonance or it is the combination of a narrow $X(4140)$ state plus a $X(4160)$ one.
@@
Among the theoretical works studying these $X$ states, some groups identified the $X(4140)$ as a molecular $D_{s}^{*}\bar{D}_{s}^{*}$ bound state with quantum numbers $0^{++}$ and $2^{++}$ [57, 58, 59]. These studies did not take into account the coupling to light meson states, thus producing a small width for the state, hence associating it to the $X(4140)$. The contribution of the light meson channels were included in Ref. [60], which generated a $2^{++}$ $X$ state at $4169~{}{\rm MeV}$ with a width of $139~{}{\rm MeV}$, coupling strongly to to $D_{s}^{*}\bar{D}_{s}^{*}$, and being associated to the $X(4160)$. In the end, the quantum numbers of the $X(4140)$ were measured, obtaining $J^{PC}=1^{++}$ [7], therefore the models that predict this state as a $D_{s}^{*}\bar{D}_{s}^{*}$ molecule are no longer supported.
Focusing on the $\Xi_{b}\to\Xi X^{*}\to\Xi~{}J/\psi~{}\phi$ decay in the present work, apart from employing it to search for the double strangeness pentaquark, we also have a chance to study the interplay between the $X(4140)$ and $X(4160)$ resonances. Similarly to what was done in Refs. [61, 62], we will use two independent primary decay mechanisms, one for the $X(4160)$ and another for the $X(4140)$. Since the $X(4140)$ resonance is not generated by any meson-meson interaction model, we assume for it a Breit-Wigner shape whose parameters are fitted to experimental data. In contrast, for the $X(4160)$ we will employ the results of a molecular model [60], which reproduces rather well its mass and width and also predicts a strong coupling to $D_{s}^{*}\bar{D}_{s}^{*}$ channel.
@@ IV.2 Primary decay
The dominant process for the $\Xi_{b}$ decay at the quark level proceeds via the external emission mechanism depicted in Fig. 15. Since the final state involves $J/\psi\,\phi$ pairs, the $D_{s}^{*}\,\bar{D}_{s}^{*}$ lines have to be connected to form a loop, as shown in Fig. 16, allowing multiple coupled channel interactions and generating dynamically the $X(4160)$, which then decays into $J/\psi$ and $\phi$. According to Ref. [60] the $X(4160)$ couples with maximum strength to the $D_{s}^{*}\,\bar{D}_{s}^{*}$ channel, but its coupling to the $J/\psi\,\phi$ channel is also significant, what also increases the dominance of the process depicted in Fig. 16. On the other hand, if we draw a diagram which would be responsible even for a direct $J/\psi\,\phi\Xi$ decay, at the microscopic quark level this reaction proceeds via the internal conversion process, shown in Fig. 17, which is strongly penalized by color factors [63] and therefore it will be neglected in this work, similarly to what was done in Ref. [61, 62].
@@
The constant $A$ on the right side of the equation represents the strength of the $\Xi_{b}\to\Xi D_{s}^{*}~{}\bar{D}_{s}^{*}$ weak decay, shown in Fig. 15. This value can be taken as a constant due to the limited range of energies involved in the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ decay as it is argued in Ref. [64]. The $(\vec{\epsilon}_{J/\psi}\times\vec{\epsilon}_{\phi})\cdot\vec{P}_{\Xi}$ factor denotes the P-wave operator which is the minimum partial wave we need in the weak vertex to conserve the angular momentum, since the spins of the $\Xi_{b}$ and $\Xi$ are $J=1/2$ while that of the $X(4160)$ is $J=2$. Here $\vec{\epsilon}_{i}~{}(i={J/\psi}\,~{}\phi)$ represent the polarization of the vector mesons in the $J/\psi~{}\phi$ rest frame and $\vec{P}_{\Xi}$ the tree-momentum of the $\Xi$ in the same frame. The factor $G_{D_{s}^{*}\bar{D}_{s}^{*}}$ indicates the contribution of the $D_{s}^{*}\bar{D}_{s}^{*}$ loop shown in Fig. 16 and $T_{D_{s}^{*}\bar{D}_{s}^{*},J/\psi\phi}$ stands for the coupled-channel unitarized amplitude for the $D_{s}^{*}\bar{D}_{s}^{*}\to J/\psi~{}\phi$ process. Note that we divide this amplitude by the couplings constants of the $X(4160)$ to the initial and final meson-baryon states. This is done for a proper comparison with the $X(4140)$ contribution, discussed in the following.
The alternative description of the decay process via the production of the $X(4140)$ is shown in Fig. 18 and the associated amplitude reads:
@@
where the $X(4140)$ resonance is parameterized with a Breit-Wigner shape, with a mass $M_{X(4140)}$ and a width $\Gamma_{X(4140)}$, and $M_{J/\psi\phi}$ is the invariant mass of the $J/\psi\phi$ system. Note the appearance of a new constant $B$, connected to the strength of the $\Xi_{b}\to\Xi~{}X(4140)$ weak decay. As we want this constant to have the same units as $A$, we introduce a $2M_{X(4140)}$ factor in the denominator. In this process we do not need to have a P-vertex since the quantum number of the $X(4140)$ is $J=1$. Therefore the minimum partial wave needed to conserve angular momentum is $L=0$.
@@ IV.3 Final state interaction
Once we discussed the generation of the $\Xi~{}J/\psi~{}\phi$ final state in the decay of the $\Xi_{b}$, we want to focus on studying the final state interaction of the $\Xi~{}J/\psi$ and $\Xi~{}\phi$ pairs.
@@
The final state interaction of $\Xi~{}J/\psi$ pairs involves connecting the $\Xi$ and $J/\psi$ legs from the diagrams of Figs. 16 and 18, forming a loop, which can dynamically generate the $S=-2$ pentaquark. The amplitude associated to the $\Xi_{b}$ decay process driven via the $X(4160)$ resonance including the final state interaction of the $J/\psi\Xi$ pair is:
@@
which is associated to the diagram shown in Fig. 19(a), where the superscript denotes the final state interaction channel. This amplitude is derived in Appendix A, where the expression for the integral $I^{J/\psi\Xi}_{X_{4160}}$ is also given. We would like to remind the reader that for the decay via an intermediate $X(4160)$ to occur the weak vertex has to be in P-wave. Therefore the operator assigned to the the $\Xi_{b}\to D^{*}_{s}~{}\bar{D}^{*}_{s}~{}\Xi$ is proportional to the momentum of the $\Xi$ and the loop integral (Fig. 19(a)) becomes a tree-vector. We evaluate this loop integral in the $J/\psi~{}\Xi$ rest frame, where the loop integral becomes proportional to $\frac{(\vec{P}_{\Xi}-\vec{P}_{\phi})}{2}$, with the scalar coefficient $I^{J/\psi\Xi}_{X_{4160}}$. The term $T_{J/\psi\Xi,J/\psi\Xi}$, which captures the final state interaction contribution from the $J/\psi~{}\Xi$ pairs, is modeled using a Breit-Wigner form:
@@
were $g_{J/\psi~{}\Xi}$ is the coupling of the $S=-2$ pentaquark to the $J/\psi~{}\Xi$ channel. $M_{J/\psi~{}\Xi}$ denotes the invariant mass of the $J/\psi~{}\Xi$ system, and $M$ and $\Gamma$ denote the mass and the width of the $S=-2$ pentaquark. The values of $g_{J/\psi~{}\Xi}$, $M$ and $\Gamma$ are taken from the molecular model of the VB interacting model discussed in Section II.
We use Breit-Wigner formula, eq. (25), instead of the full $T_{J/\psi\Xi,J/\psi\Xi}$ amplitude calculated in the previous section, because this form allows us to vary the position and the width of the pentaquark resonance in an easy and straightforward way, and, thus, to perform a more complete study on the possibility to observe double strangeness pentaquark in the $J/\psi~{}\Xi$ mass distribution from the $\Xi_{b}\to J/\psi~{}\phi~{}\Xi$ decay.
@@
where the $T_{\phi\Xi,\phi\Xi}$ amplitude is also modeled with a Breit-Wigner form, as in Eq. (25), but in this case $M$ and $\Gamma$ are the experimental mass and width of the $\Xi(2500)$. Since we do not have any model which can give us the value of $g_{\phi\Xi}$, we will invert the problem and will study for which values of $g_{\phi\Xi}$ it is feasible to see this state using the $\Xi_{b}\to\Xi~{}J/\psi~{}\phi$ reaction.
Now we go to the final state interaction in the presence of the $X(4140)$ resonance. The corresponding diagrams are given in Figs. 20(a) and 20(b) and the associated analytical expressions are:
@@
The derivation of these expressions are explained in more detail in Appendix A. The terms $I^{J/\psi\Xi}_{X_{4140}}$ and $I^{\phi\Xi}_{X_{4140}}$ are scalar loop integral analogous to those in Eqs. (24) and (26).
@@ IV.4 Full amplitude and decay rate
In this section we first proceed to combine the various terms to form the invariant amplitude in the cases where the decay driven by the $X(4140)$, ${\cal M}_{X_{4140}}$, or by the $X(4160)$, ${\cal M}_{X_{41460}}$ :
@@
where we wrote explicitly the primary decay strengths $A$ and $B$. Secondly, we add these terms obtaining the full amplitude denoted as ${\cal M}$,
@@
where the bar represents the sum over polarizations and $\beta=|B|^{2}/|A|^{2}$. Note that, since the weak decay goes in P-wave for the $X(4160)$ contribution and in S-wave for the $X(4140)$ one, the cross term in $\overline{{|\cal M}_{4160}+{\cal M}_{4140}|^{2}}$ cancels as these two partial waves are orthogonal and do not interfere. It is also important to comment that, although the the overall factor $|A|^{2}$ is not known, it is not relevant for the shape of the obtained distributions. The form of the distributions, and not its absolute value, will be our main observable. We are exploring if the position of the peak and its width can be measured in several model situations. So our final results will be given in arbitrary units and therefore we will set $|A|^{2}=1$ from here to the end of this work:
@@
The constant $\beta$ establishes a relative weight between the $X(4140)$ and $X(4160)$ contributions. Its value will be determined by following the procedure described in Ref. [62], where the interplay between the $X(4160)$ and the $X(4140)$ in the decay $B^{+}\to J/\psi~{}\phi~{}K^{+}$ is studied. The mechanism involving the $X(4160)$ in this reaction is rather similar at the quark level to our diagram shown in Fig. 15, but without the spectator $s$ quark in the initial and final state. The topology of the diagrams is also similar in the case of $X(4140)$. Therefore the $\beta$ value in our reaction may be similar as that in Ref. [62], although some difference may arise since the partial waves involved in their weak decay vertex are P-wave and D-wave for the $X(4140)$ and $X(4160)$ resonances, respectively, compared to the S- and P-waves of the present work.
We next deal with the sum over the polarizations represented by the bar in Eqs. (31) and (32). For the ${\cal M}_{X(4140)}$ amplitude this calculation is trivial and it only produces a constant, which can be reabsorbed into $\beta$. For the ${\cal M}_{X(4160)}$ amplitude the calculation is more complicated adding cross term contributions. First of all, we introduce the following definitions:
@@
Finally, the double differential cross-section for the $\Xi_{b}\to J/\Psi\phi\Xi$ decay process reads [65]:
@@
where $\mathcal{M}$ is given by Eq. (32). We wrote in Eq. (IV.4) the double differential cross section as a function of the $M_{\phi\Xi}$ and $M_{J/\psi\Xi}$ invariant masses, but we could have used any other combination of the two particle invariant masses of the $J/\Psi\phi\Xi$ final state.
Now, fixing the invariant mass $M_{J/\psi\Xi}$, one can integrate over $M_{\phi\Xi}$ in order to obtain $d\Gamma/dM_{J/\psi\Xi}$. In this case, the integration limits are given by: